OpenStudy (barreraa):

Use the quadratic equation y = x2 + 3x + 9. Where does the vertex of the parabola lie? (Points : 5) below the x-axis above the x-axis on the x-axis on the y-axis

OpenStudy (barreraa):

i think this one is A @sasogeek

Nnesha (nnesha):

\[\huge\rm x=\frac{ -b }{ 2a }\] formula to find x coordinate of vetex

OpenStudy (barreraa):

ok thanks but i just want to know if im right or wrong @Nnesha

Nnesha (nnesha):

nope that's incorrect substitute x value y=x^2+3x+9 solve for y y is positive or negative ?

OpenStudy (barreraa):

i think its between B and D @Nnesha

Nnesha (nnesha):

so which one do not guess :-)

Nnesha (nnesha):

\[\huge\rm Ax^2+Bx+C=0\] quadratic equation A leading coefficient C is constant term

OpenStudy (barreraa):

b above the x-axis @Nnesha

Nnesha (nnesha):

is that an educated guess or are u 100% sure ?? :=)

OpenStudy (barreraa):

thanks for the help @Nnesha

Nnesha (nnesha):

np :=)